A boat is floating at rest in a dense fog near a large cliff. The captain sounds its horn at water level and the sound travels through the salt water (1470 m/s) and the air (340m/s). The echo in the water (due to the cliff) takes 0.40s to return. How much additional time will it take for the echo in the air to travel?
1
Expert's answer
2020-12-29T15:32:16-0500

"l/1470=0.4\to l=588(m)"


"588\/340=1.73(s)"


"1.73-0.4=1.33(s)" . Answer